# Spherical maximal functions and Hardy spaces for Fourier integral operators

### v2

@@ 3.2. Hypersurfaces
It is natural to ask whether one can extend (3.13) to more general $f\in\mathcal{H}^{s,p}_{FIO}(\mathbb{R}^{n})$. Here one approach is to recall that, for $f\in C(\mathbb{R}^{n})$, one has $\mathcal{A}_{t}f=b(tD)f$ for $b=\widehat{\psi\mathrm{d}\sigma}$. Since $b(tD):\mathcal{S}^{\prime}(\mathbb{R}^{n})\to\mathcal{S}^{\prime}(\mathbb{R}^{n})$ continuously, $\mathcal{A}_{t}f$ can be extended in a distributional sense to general $f\in\mathcal{S}^{\prime}(\mathbb{R}^{n})$. Moreover, $b(tD)f\in L^{p}(\mathbb{R}^{n})$ if $f\in\mathcal{H}^{s,p}_{FIO}(\mathbb{R}^{n})$, as follows from Proposition 2.2 and the proof of Theorem 3.6, so that this extension of (3.12) also makes sense pointwise almost everywhere. Of course, there is a subtlety here regarding the choice of a pointwise representative of $b(tD)f\in L^{p}(\mathbb{R}^{n})$ for different values of $t\in[1,2]$, but this can be dealt with as in the proof of Theorem 3.1. By approximation, (3.13) then extends to all $f\in\mathcal{H}^{s,p}_{FIO}(\mathbb{R}^{n})$ if $p<\infty$.
On the other hand, one might wonder whether this extension of $\mathcal{M}_{\Sigma,\psi}f$ can again be expressed using integrals as in (3.11). By the reasoning in [36, Section XI.3.5], this is indeed the case if, for example, $p<\infty$ and $f\in L^{\infty}_{\mathrm{loc}}(\mathbb{R}^{n})\cap\mathcal{H}^{s,p}_{FIO}(\mathbb{R}^{n})$. On the other hand, if $p>\frac{2}{n-1}$ and $s\in(d(p)+\frac{1}{p}-\frac{n-1}{2},s(p))$, then there exist $f\in\mathcal{H}^{s,p}_{FIO}(\mathbb{R}^{n})$ which are not contained in $L^{p}_{\mathrm{loc}}(\mathbb{R}^{n})$. For such $f$ it is less clear how to make sense of the integral in (3.11).
A similar remark applies to Theorem 3.9 and Proposition 3.15 below.
@@
    has $n-1$ non-vanishing principal curvatures.
@@
of $\{(x,t,y)\mid(x,t)\in\mathbb{R}^{n+1},y\in\Sigma_{x,t}\}$ satisfies the cinematic curvature condition from [32] (see also [33, Section 8.1] and Remark 2.4).
Let $\mathrm{d}\sigma_{x,t}$ be Lebesgue measure restricted to $\Sigma_{x,t}$, and let $\rho\in C^{\infty}_{c}(\mathbb{R}^{n+1}\times\mathbb{R}^{n})$. For $f\in C(\mathbb{R}^{n})$, $x\in\mathbb{R}^{n}$ and $t\in\mathbb{R}$, set
@@
We then obtain a version of Theorem 3.6 for varying hypersurfaces. By Theorem 4.9 (see also Remark 4.10), this result is essentially sharp for all $p\in[1,2]\cup[\frac{2(n+1)}{n-1},\infty]$ whenever $\rho$ is not identically zero.
@@
Let $\Psi\in C^{\infty}(\mathbb{R}^{n+1}\times\mathbb{R}^{n})$ be such that (1) and (2) hold, and let $\rho\in C^{\infty}_{c}(\mathbb{R}^{n+1}\times\mathbb{R}^{n})$. For $t\in\mathbb{R}$, let $T_{t}$ be as in (3.17). Then, for all $p\in[1,\infty]$ and $s>d(p)+\frac{1}{p}-\frac{n-1}{2}$, there exists a $C\geq 0$ such that
@@
We will reduce to the setting of Theorem 3.4 by decomposing into finitely many operators and applying changes of coordinates. However, some care has to be taken to deal with the fact that (3.18) is not invariant under general changes of coordinates in the $(x,t)$ variable in (3.17).
@@
for all $f\in C(\mathbb{R}^{n})$, $t\in\mathbb{R}$ and $x\in\mathbb{R}^{n}$, where $\rho_{0}\in C^{\infty}_{c}(\mathbb{R}^{n+1}\times\mathbb{R})$ satisfies $\rho_{0}(x,t,y)=\rho(x,t,y)|\partial_{y}\Psi_{t}(x,y)|$ for $y\in\Sigma_{x,t}$ (see e.g. [36, Section XI.3.1.2]). Hence the kernel of $T_{t}$ is a Lagrangian distribution of order $-(n-1)/2$, and $T_{t}$ is a Fourier integral operator of order $-(n-1)/2$ with canonical relation
@@
Condition (1) implies that $\mathcal{C}_{t}$ is a local canonical graph (see e.g. [33, p. 171]).
@@
Now, by e.g. [20, Proposition A.2], there exists a collection $(T_{j})_{j=1}^{l}$ of operators as in Proposition 2.3, with $m=-(n-1)/2$, changes of coordinates $(S_{j})_{j=1}^{l}$ on $\mathbb{R}^{n}$, and an operator $R$ with a kernel which is a Schwartz function, such that $T=\sum_{j=1}^{l}T_{j}S_{j}+R$ (that no changes of coordinates on $\mathbb{R}^{n+1}$ are required is noted in [20, Remark A.3]). The smoothing operator $R$ satisfies the required maximal function estimate, since it maps $\mathcal{H}^{s,p}_{FIO}(\mathbb{R}^{n})$ into $\mathcal{S}(\mathbb{R}^{n+1})$ continuously. Moreover, $\mathcal{H}^{s,p}_{FIO}(\mathbb{R}^{n})$ is invariant under the changes of coordinates $S_{j}$, and each $T_{j}$ satisfies the required maximal function estimate, by Theorem 3.4. This concludes the proof. ∎
@@
for all $p\in[1,\infty]$, $s>s(p)+\frac{1}{p}-\frac{n-1}{2}$ and $f\in C(\mathbb{R}^{n})\cap\mathcal{H}^{s,p}_{FIO}(\mathbb{R}^{n})$. This follows as before, if one relies on Remark 3.5 instead of Theorem 3.4.
@@
for all $p\in[1,\infty]$, $s>s(p)+\frac{1}{p}-\frac{n-1}{2}$ and $f\in C(\mathbb{R}^{n})\cap\mathcal{H}^{s,p}_{FIO}(\mathbb{R}^{n})$. This follows as in the proof of Theorem 3.6, but relying on (3.10) instead of Theorem 3.1.
@@ 3.3. Spherical means of complex order
for all $\xi\neq 0$, where $J_{\beta}$ is the Bessel function of the first kind, of order $\beta\in\mathbb{C}$ (see [37, Theorem IV.4.15]). For a general $\alpha\in\mathbb{C}$ one can then also define $m_{\alpha}$ as in (3.21). In turn, this then leads to the definition of the complex spherical mean $\mathcal{M}^{\alpha}_{t}$, by setting
@@
$\big{\|}\sup_{t\in I}|\mathcal{M}^{\alpha}_{t}f|\big{\|}_{L^{p}(\mathbb{R}^{n})}\leq C\|f\|_{\mathcal{H}^{s,p}_{FIO}(\mathbb{R}^{n})}$
@@
Note that, as follows from asymptotics of Bessel functions (see [38, Chapters 3 and 7]), there exist $a^{0}_{\alpha}\in C^{\infty}_{c}(\mathbb{R}^{n})$ and $a^{+}_{\alpha},a^{-}_{\alpha}\in S^{-\frac{n-1}{2}-\operatorname{Re}(\alpha)}(\mathbb{R}^{n})$ such that
@@
Now, as in the proof of Theorem 3.1, for all $f\in\mathcal{S}(\mathbb{R}^{n})$ one has
$\big{\|}\sup_{t\in I}|a^{0}_{\alpha}(tD)f|\big{\|}_{L^{p}(\mathbb{R}^{n})}\lesssim\|f\|_{\mathcal{H}^{s,p}_{FIO}(\mathbb{R}^{n})}.$
Moreover, Theorem 3.1 yields the required bounds for the remaining terms in (3.23). ∎
@@
The expression for $\mathcal{M}^{\alpha}_{t}f$ in (3.20), for $\alpha>0$ and $f\in C(\mathbb{R}^{n})$, extends to $\alpha\in\mathbb{C}$ with $\operatorname{Re}(\alpha)>0$. Hence, if $f\in C(\mathbb{R}^{n})\cap\mathcal{H}^{s,p}_{FIO}(\mathbb{R}^{n})$, then Proposition 3.11 yields maximal function bounds for these integrals. As in Remark 3.7, these estimates can in turn be extended to $f\in L^{\infty}_{\mathrm{loc}}(\mathbb{R}^{n})\cap\mathcal{H}^{s,p}_{FIO}(\mathbb{R}^{n})$. Moreover, the bounds in this subsection extend to $p=\infty$, as in Remark 3.3.
@@ 3.4. Maximal functions on manifolds
Now, for $f\in\mathcal{D}(M,\Omega_{1/2})$, $t\in\mathbb{R}$ and $x\in M$, set $Tf(x,t):=e^{it\sqrt{-\Delta_{g}}}f(x)$. By e.g. [17, Theorem 29.1.1 and Proposition 29.1.9], $T$ is an FIO of order $-1/4$ from $M$ to $M\times\mathbb{R}$, with canonical relation
@@
Here $(\chi_{t})_{t\in\mathbb{R}}$ is the geodesic flow on $T^{*}M$. In particular, each $e^{it\sqrt{-\Delta_{g}}}$ is a Fourier integral operator of order 00, the canonical relation of which is the graph of $\chi_{t}$. This in turn means that $e^{it\sqrt{-\Delta_{g}}}$ acts continuously on $\mathcal{D}(M,\Omega_{1/2})$ and $\mathcal{D}^{\prime}(M,\Omega_{1/2})$, and $e^{it\sqrt{-\Delta_{g}}}f\in\mathcal{D}^{\prime}(M,\Omega_{1/2})$ is a priori well defined for each $f\in\mathcal{H}^{s,p}_{FIO}(M)$. Moreover, if $f\in\mathcal{D}(M,\Omega_{1/2})$ then the left-hand side of (3.25) is well defined.
Finally, $\mathcal{C}$ satisfies the cinematic curvature condition (see e.g. [33, Section 8.1]). Thus, by working in local coordinates, one can argue as in the proof of Theorem 3.9 to derive the required statement from Theorem 3.4. ∎
@@
for $t\in(0,\mathrm{inj}(M))$, $f\in C(M,\Omega_{1/2})$ and $x\in M$.
@@
Let $p\in[1,\infty]$ and $s>d(p)+\frac{1}{p}-\frac{n-1}{2}$. Then, for each compact interval $I\subseteq(0,\mathrm{inj}(M))$, there exists a $C\geq 0$ such that
@@
If one sets $Tf(x,t):=\mathcal{A}_{t,g}f(x)$ for $f\in\mathcal{D}(M,\Omega_{1/2})$ and $(x,t)\in M\times I$, then the canonical relation of $T$ is the union of the relation in (3.27) and the corresponding one with the $\tau$ coordinate replaced by $-\tau$. Since each $\mathcal{A}_{t,g}$ is an FIO of order $-(n-1)/2$, one can now argue as in the proof of Proposition 3.14.
Alternatively, one can reason as follows. It suffices to work in a sufficiently small geodesic normal coordinate neighborhood of a given point $x_{0}\in M$. There each $\mathcal{A}_{t,g}$ is as in (3.17), with $\Sigma_{x_{0},t}=tS^{n-1}$ and $\Psi_{t}(x_{0},y)=|x_{0}-y|-t$. Then $\Psi_{t}(x,y)$ is a small perturbation of $\tilde{\Psi}_{t}(x,y):=|x-y|-t$ for $x$ near $x_{0}$. Since $\tilde{\Psi}_{t}$ satisfies conditions (1) and (2), and because these conditions are stable under small perturbations, the required statement follows from Theorem 3.9. ∎

### v3

@@ 3.2. Hypersurfaces
It is natural to ask whether one can extend (3.13) to more general $f\in\mathcal{H}^{s,p}_{FIO}(\mathbb{R}^{n})$. Here one approach is to recall that, if $f\in C(\mathbb{R}^{n})\cap\mathcal{H}^{s,p}_{FIO}(\mathbb{R}^{n})$, then $\mathcal{A}_{t}f=b(tD)f$ as distributions, where $b=\widehat{\psi\mathrm{d}\sigma}$. Moreover, as in the proof of Theorem 3.7 one can show that, by identifying $\mathcal{A}_{t}f$ and $b(tD)f$, one obtains an extension of (3.12) to general $f\in\mathcal{H}^{s,p}_{FIO}(\mathbb{R}^{n})$ which makes sense pointwise almost everywhere. Of course, there is a subtlety here regarding the choice of a pointwise representative of $b(tD)f\in L^{p}(\mathbb{R}^{n})$ for different values of $t\in[1,2]$, but this can be dealt with in the exact same manner as in the proof of Theorem 3.2. More precisely, one can choose these representatives in such a way that $x\mapsto\sup_{1\leq t\leq 2}|b(tD)f(x)|$ is measurable. By approximation, (3.13) then extends to all $f\in\mathcal{H}^{s,p}_{FIO}(\mathbb{R}^{n})$.
Now, one might wonder whether the extension of (3.13) to general $f\in\mathcal{H}^{s,p}_{FIO}(\mathbb{R}^{n})$ using $b(tD)$ can again be expressed in terms of integrals as in (3.11). By the reasoning in [40, Section XI.3.5], this is indeed the case if $p>\frac{n}{n-1}$ and $f\in L^{p}(\mathbb{R}^{n})$, given that then $L^{p}(\mathbb{R}^{n})\subseteq\mathcal{H}^{s,p}_{FIO}(\mathbb{R}^{n})$ continuously for $s\in(d(p)+\frac{1}{p}-\frac{n-1}{2},-s(p)]$. On the other hand, if $p>\frac{2}{n-1}$ and $s\in(d(p)+\frac{1}{p}-\frac{n-1}{2},s(p))$, then $\mathcal{H}^{s,p}_{FIO}(\mathbb{R}^{n})$ is not a subset of $L^{p}(\mathbb{R}^{n})$, by the sharpness of the embeddings in (2.2). In fact, in this case there exist $f\in\mathcal{H}^{s,p}_{FIO}(\mathbb{R}^{n})$ which are not contained in $L^{p}_{\mathrm{loc}}(\mathbb{R}^{n})$; one such example can essentially be found in [40, Section IX.6.13]. For such $f$ it is less clear how to make sense of the integral in (3.11).
Similar comments apply to Theorem 3.10 and Proposition 3.16 below.
@@
    has n1n-1 non-vanishing principal curvatures.
@@
of $\{(x,t,y)\mid(x,t)\in\mathbb{R}^{n+1},y\in\Sigma_{x,t}\}$ satisfies the cinematic curvature condition from [36] (see also [37, Section 8.1] and Remark 2.4).
Let $\mathrm{d}\sigma_{x,t}$ be Lebesgue measure restricted to $\Sigma_{x,t}$, defined as before, and let $\rho\in C^{\infty}_{c}(\mathbb{R}^{n+1}\times\mathbb{R}^{n})$. For $f\in C(\mathbb{R}^{n})$, $x\in\mathbb{R}^{n}$ and $t\in\mathbb{R}$, set
@@
We then obtain a version of Theorem 3.7 for varying hypersurfaces. By Theorem 4.9 (see also Remark 4.10), this result is essentially sharp for all $p\in[1,2]\cup[\frac{2(n+1)}{n-1},\infty)$ whenever $\rho$ is not identically zero.
@@
Let $\Psi\in C^{\infty}(\mathbb{R}^{n+1}\times\mathbb{R}^{n})$ be such that (1) and (2) hold, and let $\rho\in C^{\infty}_{c}(\mathbb{R}^{n+1}\times\mathbb{R}^{n})$. For $t\in\mathbb{R}$, let $T_{t}$ be as in (3.20). Then, for all $p\in[1,\infty)$ and $s>d(p)+\frac{1}{p}-\frac{n-1}{2}$, there exists a $C\geq 0$ such that
@@
Again, the left-hand side is well defined for general $f\in C(\mathbb{R}^{n})\cap\mathcal{H}^{s,p}_{FIO}(\mathbb{R}^{n})$. We will reduce to the setting of Theorem 3.5 by decomposing into finitely many operators and applying changes of coordinates. However, some care has to be taken to deal with the fact that (3.21) is not invariant under general changes of coordinates in the $(x,t)$ variable in (3.20).
@@
for all $f\in C(\mathbb{R}^{n})$, $t\in\mathbb{R}$ and $x\in\mathbb{R}^{n}$, where $\rho_{0}\in C^{\infty}_{c}(\mathbb{R}^{n+1}\times{\mathbb{R}^{n}})$ satisfies $\rho_{0}(x,t,y)=\rho(x,t,y)|\partial_{y}\Psi_{t}(x,y)|$ for $y\in\Sigma_{x,t}$ (see e.g. [40, Section XI.3.1.2]). Hence the kernel of $T_{t}$ is a Lagrangian distribution of order (n1)/2-(n-1)/2, and $T_{t}$ is a Fourier integral operator of order (n1)/2-(n-1)/2 with canonical relation
@@
Condition (1) implies that $\mathcal{C}_{t}$ is a local canonical graph (see e.g. [37, p. 171]).
@@
Now, by e.g. [22, Proposition A.2], there exists a collection $(T_{j})_{j=1}^{l}$ of operators as in Proposition 2.3, with m=(n1)/2m=-(n-1)/2, changes of coordinates $(S_{j})_{j=1}^{l}$ on $\mathbb{R}^{n}$, and an operator $R$ with a kernel which is a Schwartz function, such that $T=\sum_{j=1}^{l}T_{j}S_{j}+R$ (that no changes of coordinates on $\mathbb{R}^{n+1}$ are required is noted in [22, Remark A.3]). The smoothing operator $R$ satisfies the required maximal function estimate, since it maps $\mathcal{H}^{s,p}_{FIO}(\mathbb{R}^{n})$ into $\mathcal{S}(\mathbb{R}^{n+1})$ continuously. Moreover, $\mathcal{H}^{s,p}_{FIO}(\mathbb{R}^{n})$ is invariant under the changes of coordinates $S_{j}$, and each $T_{j}$ satisfies the required maximal function estimate, by Theorem 3.5. This concludes the proof. ∎
@@
for all $p\in[1,\infty)$, $s>s(p)+\frac{1}{p}-\frac{n-1}{2}$ and $f\in C(\mathbb{R}^{n})\cap\mathcal{H}^{s,p}_{FIO}(\mathbb{R}^{n})$. This follows as before, if one relies on Remark 3.6 instead of Theorem 3.5.
@@
for all $p\in[1,\infty)$, $s>s(p)+\frac{1}{p}-\frac{n-1}{2}$ and $f\in C(\mathbb{R}^{n})\cap\mathcal{H}^{s,p}_{FIO}(\mathbb{R}^{n})$. This follows as in the proof of Theorem 3.7, but relying on (3.10) instead of Theorem 3.2.
@@ 3.3. Spherical means of complex order
for all $\xi\neq 0$, where $J_{\beta}$ is the Bessel function of the first kind, of order $\beta\in\mathbb{C}$ (see [41, Theorem IV.4.15]). For a general $\alpha\in\mathbb{C}$ one can then also define $m_{\alpha}$ as in (3.24). In turn, this then leads to the definition of the complex spherical mean $\mathcal{M}^{\alpha}_{t}$, by setting
@@
(3.26) $\big\|\sup_{t\in I}|\mathcal{M}^{\alpha}_{t}f|\big\|_{L^{p}(\mathbb{R}^{n})}\leq C\|f\|_{\mathcal{H}^{s,p}_{FIO}(\mathbb{R}^{n})}$
@@
Note that, as follows from asymptotics of Bessel functions (see [43, Chapters 3 and 7]), there exist $a^{0}_{\alpha}\in\cap_{m\in{\mathbb{R}}}S^{m}(\mathbb{R}^{n})$ and $a^{+}_{\alpha},a^{-}_{\alpha}\in S^{-\frac{n-1}{2}-\operatorname{Re}(\alpha)}(\mathbb{R}^{n})$ such that
@@
Now, as in the proof of Theorem 3.2, if $f\in\mathcal{S}(\mathbb{R}^{n})$ then
$\big\|\sup_{t\in I}|a^{0}_{\alpha}(tD)f|\big\|_{L^{p}(\mathbb{R}^{n})}\lesssim\|f\|_{\mathcal{H}^{s,p}_{FIO}(\mathbb{R}^{n})}$
is a well-defined inequality which is valid. Theorem 3.2 yields the required bounds for the other terms in (3.27). Finally, as in the proof of Theorem 3.2, by approximation (3.26) extends in a well-defined manner to general $f\in\mathcal{H}^{s,p}_{FIO}(\mathbb{R}^{n})$. ∎
@@
The expression for $\mathcal{M}^{\alpha}_{t}f$ in (3.23), for $\alpha>0$ and $f\in C(\mathbb{R}^{n})$, extends to $\alpha\in\mathbb{C}$ with $\operatorname{Re}(\alpha)>0$. Hence, if $f\in C(\mathbb{R}^{n})\cap\mathcal{H}^{s,p}_{FIO}(\mathbb{R}^{n})$, then Proposition 3.12 yields maximal function bounds for these integrals. As in Remark 3.8, these estimates can in turn be extended to $f\in L^{\infty}_{\mathrm{loc}}(\mathbb{R}^{n})\cap\mathcal{H}^{s,p}_{FIO}(\mathbb{R}^{n})$.
@@ 3.4. Maximal functions on manifolds
Now, for $f\in\mathcal{D}(M,\Omega_{1/2})$, $t\in\mathbb{R}$ and $x\in M$, set $Tf(x,t):=e^{it\sqrt{-\Delta_{g}}}f(x)$. By e.g. [19, Theorem 29.1.1 and Proposition 29.1.9], $T$ is an FIO of order 1/4-1/4 from $M$ to $M\times\mathbb{R}$, with canonical relation
@@
Here $(\chi_{t})_{t\in\mathbb{R}}$ is the geodesic flow on $T^{*}M$. In particular, each $e^{it\sqrt{-\Delta_{g}}}$ is a Fourier integral operator of order 0, the canonical relation of which is the graph of $\chi_{t}$. This in turn means that $e^{it\sqrt{-\Delta_{g}}}$ acts continuously on $\mathcal{D}(M,\Omega_{1/2})$ and $\mathcal{D}^{\prime}(M,\Omega_{1/2})$, and $e^{it\sqrt{-\Delta_{g}}}f\in\mathcal{D}^{\prime}(M,\Omega_{1/2})$ is a priori well defined for each $f\in\mathcal{H}^{s,p}_{FIO}(M)$. Moreover, if $f\in\mathcal{D}(M,\Omega_{1/2})$ then the left-hand side of (3.29) is well defined.
Finally, $\mathcal{C}$ satisfies the cinematic curvature condition (see e.g. [37, Section 8.1]). Thus, by working in local coordinates, one can argue as in the proof of Theorem 3.10 to derive the required statement from Theorem 3.5. ∎
@@
for $t\in(0,\mathrm{inj}(M))$, $f\in C(M,\Omega_{1/2})$ and $x\in M$, where we identify $\sqrt{dV_{g}}$ with its restriction to $S_{t}(x)$, defined in an analogous manner as in Section 3.2.
@@
Let $p\in[1,\infty)$ and $s>d(p)+\frac{1}{p}-\frac{n-1}{2}$. Then, for each compact interval $I\subseteq(0,\mathrm{inj}(M))$, there exists a $C\geq 0$ such that
@@
If one sets $Tf(x,t):=\mathcal{A}_{t,g}f(x)$ for $f\in\mathcal{D}(M,\Omega_{1/2})$ and $(x,t)\in M\times I$, then the canonical relation of $T$ is the union of the relation in (3.31) and the corresponding one with the $\tau$ coordinate replaced by $\tau$. Since each $\mathcal{A}_{t,g}$ is an FIO of order (n1)/2-(n-1)/2, one can now argue as in the proof of Proposition 3.15.
Alternatively, one can reason as follows. It suffices to work in a sufficiently small geodesic normal coordinate neighborhood of a given point $x_{0}\in M$. There each $\mathcal{A}_{t,g}$ is as in (3.20), with $\Sigma_{x_{0},t}=x_{0}+tS^{n-1}$ and $\Psi_{t}(x_{0},y)=|x_{0}-y|-t$. Then $\Psi_{t}(x,y)$ is a small perturbation of $\tilde{\Psi}_{t}(x,y):=|x-y|-t$ for $x$ near $x_{0}$. Since $\tilde{\Psi}_{t}$ satisfies conditions (1) and (2), and because these conditions are stable under small perturbations, the required statement follows from Theorem 3.10. ∎